# VIALM: A Survey and Benchmark of Visually Impaired Assistance with Large Models

### 29 Jan 2024 v1

@@ 3.1 Evaluation Dataset
We collected images from the Internet of the two most common daily life environments: indoor homes and supermarkets. The evaluation dataset was subsequently manually crafted by employing highly educated annotators with English-medium undergraduate or graduate degrees (Li et al., 2023b). Figure 3 illustrates a test data sample.
Dataset Statistics
The evaluation dataset consists of 200 samples in total, divided into two scenarios with 100 images each, similar in size to Chen et al. (2023b). Every image is paired with a corresponding question and answer. The questions are designed to prompt the models to offer guidance for VI users, utilizing tactile clues to navigate towards specific targets. Notably, the dataset features a broad coverage, with 96 different target objects, such as ’meat’ and ’basin’ in the supermarket scenario, and ’sofa’ and ’fridge’ in the home scenario. Furthermore, in comparison to benchmark datasets like OK-VQA (average answer length: 8.1 tokens) (Marino et al., 2019) and CLEVR (average answer length: 18.4 tokens) (Johnson et al., 2017), our dataset exhibits longer questions (average question length: 51 tokens) and significantly longer answers (average answer length: 99.8 tokens). It indicates a higher level of difficulty for models in question answering.
@@ 3.2 Evaluation Metrics
Automatic Evaluation Metrics
@@
Human Evaluation Metrics
The human evaluation measures the models’ performance in three main aspects: Correctness, Actionability, and Clarity, culminating in an overall helpfulness score. Correctness evaluates the accuracy of the guidance in describing the environment and indicating target positions. Actionability evaluates the user’s ease in following the provided guidance. Clarity assesses the comprehensibility of the answer’s language from the user’s perspective. The overall helpfulness score is calculated using weights of 0.5 for Actionability, 0.4 for Correctness, and 0.1 for Clarity.
@@ 4 Experiment Results
Figure 4: Evaluation Results. (a) Overall average experimental results for six models—GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA, BLIVA—based on ROUGE, BERTScore, and Human Evaluation metrics. (b) Human evaluation results of the six models focusing on the aspects of Correctness, Actionability, and Clarity.
We present the overall experimental results in Figure 4 (a). First, GPT-4 and LLaVA both exhibit relatively high BERTSCore values (above 0.6 within a range of 0 to 1), indicating that current SOTA VLMs possess the potential for the VIALM task (Viability). Second, regarding ROUGE, all models demonstrate similar but low values (below 0.2 within a range of 0 to 1), suggesting that the model predictions may be further improved. Third, the human evaluation results, which have been scaled into a range from 0 to 1, reveal that GPT-4 performs the best, with LLaVA and CogVLM following as the second-best performers, motivating further study.
@@ 4.1 Automatic Evaluation Results
Table 4: Automatic Evaluation Results. Specific versions are GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2-7B, Qwen-VL-Chat-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B. Except for GPT-4, the others operated with float16 precision. The highest and runner-up metric values are distinctly marked in gray. ’RG’ represents ROUGE, ’BERT_S’ stands for BERTScore, and ’Len’ signifies output length.
Table 4 displays automatic evaluation results for the two distinct scenarios. In both situations,GPT-4 and LLaVA possess the highest BERTScores,indicating their superior abilities. Additionally, when considering ROUGE, we can see: (1) all models perform better in the home scenario setting; (2) LLaVA displays relatively high ROUGE values and CogVLM excels in a more challenging environment; (3) GPT-4 performs worse than smaller-scaled models. The first observation suggests that the increased visual complexity of the supermarket environment (i.e., a greater variety of objects and more colors in the image) makes the VIALM task more challenging, highlighting the need for models to generalize with environment adaptability. Combining the second and third observations and analyzing the generated guidance, we identify GPT-4’s significant drawback: its output is excessively redundant and lengthy, resulting in low ROUGE scores and reduced functionality. In contrast, CogVLM generates the shortest outputs, while LLaVA exhibits the best performance with a moderate guidance length and high ROUGE scores.
4.2 Human Evaluation Results
@@
Table 5: Human Evaluation Fail Rates. ’Not Grounded Rate’ is calculated by the sample proportion with Correctness $\leq$ 3 (Score range: 1–5). ’Not Fine-grained Rate’ is derived from Actionability $\leq$ 3.
Figure 4 (b) presents the human evaluation results. To mitigate bias, two independent participants conducted evaluations. In each distinct scenario, 50 random samples were selected. Pearson and Spearman correlation coefficients between the two annotators for each dimension range from 0.7 to 0.9, indicating a strong consensus on the models’ performance. First, it is once again observed that all models provide more accurate answers in home settings than in supermarket settings. Second, as shown in Figure 4 (b), GPT-4 excels in performance across all three dimensions, with CogVLM and LLaVA closely following in second place. Subsequently, we analysis the results for the three specific dimensions:
$\bullet$ In Correctness, GPT-4 performs the best, with CogVLM slightly surpassing LLaVA. It might indicate that CogVLM understands the visual environment better than LLaVA, and therefore leads to a superior environment grounding ability.
$\bullet$ In Actionability, GPT-4 performs the best, with LLaVA narrowly outperforming CogVLM. Since we prompted the use of tactile clues, this observation implies LLaVA’s LLM component, LLaMA-2, generates more effective guidance to VI users, incorporating tactile and fine-grained instructions.
@@
This analysis reveals two prevalent shortcomings shared by all six models, summarized as ’not grounded’(indicating low correctness) and ’not fine-grained’ (indicating low actionability). The failure rates are presented in Table 5. It is notable that, despite GPT-4’s superior performance, 25.7% of its responses are not grounded and 32.1% not fine-grained.
4.3 Case Study
To demonstrate the models’ output, we present an example case in Figure 5 where the question seeks guidance to obtain mustard. The predictions of GPT-4, CogVLM, and LLaVA are provided, yielding the following observations: (1) GPT-4 generates numerous tokens offering general guidance but lacks image grounding. In addition, its guidance is not easy for VI users to follow. (2) CogVLM exhibits better environment grounding, specifying the object’s position to the user’s left, yet it lacks detailed step-by-step guidance. (3) LLaVA’s output emphasizes guiding the user to the specific object but appears to have limited understanding of the environment.
@@ 4.4 Discussions
Model Rouge_1 Rouge_2 Rouge_L BERTScore
Grounded Part
GPT-4 0.156 0.028 0.104 0.405
CogVLM 0.262 0.043 0.158 0.284
MiniGPT4 0.211 0.053 0.139 0.415
Qwen-VL 0.201 0.028 0.132 0.311
LLaVA 0.186 0.039 0.123 0.404
BLIVA 0.133 0.025 0.091 0.375
Fine-Grained Part
GPT-4 0.292 0.072 0.166 0.573
CogVLM 0.317 0.083 0.189 0.443
MiniGPT4 0.295 0.062 0.177 0.455
Qwen-VL 0.331 0.088 0.193 0.506
LLaVA 0.341 0.094 0.202 0.544
BLIVA 0.259 0.062 0.159 0.514
@@
This section explores model selection strategies to improve VIALM, particularly in addressing the two mentioned limitations. We divided the ground truth text into two parts: the grounded part (i.e., environmental), and the fine-grained part (i.e., tactile). We then compared model predictions with each part and calculated automatic metric scores. Since Table 4 reveals that Qwen-VL and LLaVA exhibit superior performance under the home scenario, a notable deviation from the supermarket scenario and our human evaluation results, we focus on the home scenario. Results are displayed in Table 6.
Analyzing Table 6 reveals two significant findings. Firstly, CogVLM excels in environment grounding, possibly due to two factors: (1) the incorporation of the more recent Eva-Clip ViT model, enhancing its visual perception capabilities, and (2) the inclusion of a Visual Expert module in CogVLM’s LLM architecture, augmenting its visual-language alignment abilities. Secondly, LLaVA demonstrates superior efficiency in generating fine-grained guidance, possibly because its LLM component, LLaMA-2, might possess advanced capabilities concerning tactile description. These initial findings can assist researchers in developing VIALM models.
@@ 5 Conclusion
In this study, we defined the VIALM task and carried out a comprehensive survey of LMs for VIA, encompassing LLMS, VLMs, and embodied agents. Furthermore, we constructed a benchmark dataset with comprehensive coverage and high difficulty. Subsequently, we conducted a benchmark evaluation to measure the zero-shot performance of LMs, aiming to examine the potential and limitations of current SOTA models in the VIALM task. The experimental results revealed two primary limitations in the models: their ability to generate grounded and fine-grained guidance. To address these limitations, the experiment results suggest that improving visual processing can aid in producing more grounded advice, and incorporating the tactile modality can contribute to generating more fine-grained guidance for VI users.

### 10 Feb 2024 v2

@@ 3.1 Evaluation Dataset
We collected images from the Internet of the two most common daily life environments: homes and supermarkets. The dataset was subsequently manually curated by employing well-educated and English-fluent annotators, following the practice in Li et al. (2023b). Specifically, the ground truth construction is supervised by VIA experts at The Hong Kong Society for the Blind (HKSB).
Dataset Characteristics.
Each image in the dataset is paired with a corresponding question and an answer. Questions are crafted to encourage the models to offer guidance (answers) tailored for VI users. In formulating answers, tactile sensation is considered, as previous work indicates that tactile perception is more pronounced among the VI group, compensating for vision loss during task completion (Ottink et al., 2022). Figure 3 shows an example from the dataset.
Dataset Statistics.
The evaluation dataset consists of 200 samples in total, divided into two scenarios with 100 images each, similar in size to Chen et al. (2023b). It features a broad coverage, with 96 different target objects, such as ‘meat’ and ‘basin’ in the supermarket scenario, and ‘sofa’ and ‘fridge’ in the home scenario. Furthermore, in comparison to benchmark datasets like OK-VQA (average answer length: 8.1 tokens) (Marino et al., 2019) and CLEVR (average answer length: 18.4 tokens) (Johnson et al., 2017), our dataset exhibits longer questions (average question length: 51 tokens) and significantly longer answers (average answer length: 99.8 tokens). It indicates a higher level of difficulty for models.
@@ 3.2 Evaluation Metrics
Automatic Evaluation Metrics.
@@
Human Evaluation Metrics.
[…]
@@ 4 Experiment Results
Figure 4: Evaluation Results. (a) The overall average automatic evaluation results for these six models: GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA, and BLIVA, based on ROUGE and BERTScore metrics. (b) The human evaluation results of the six models focusing on the aspects of Correctness, Actionability, and Clarity.
@@ 4.1 Automatic Evaluation Results
Table 4: Automatic Evaluation Results. Specific versions are GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2-7B, Qwen-VL-Chat-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B. Except for GPT-4, the others operated with float16 precision. The highest and runner-up metric values are distinctly marked in gray. ‘RG’ represents ROUGE, ‘BERT_S’ stands for BERTScore, and ‘Len’ signifies output length.
The overall automatic experimental results are illustrated in Figure 4 (a). We observe GPT-4 and LLaVA exhibit relatively high BERTSCore values (above 0.6 within a range of 0 to 1). For ROUGE, all models demonstrate consistently low values (below 0.2 within a range of 0 to 1). Given that BERTScore quantifies semantic similarity while ROUGE measures strict matching to the references, this observation illustrates that LMs can generally address the task, though without strict matching, indicating LMs’ fundamental VIA skills (RQ1).
In a detailed analysis, Table 4 presents automatic evaluation results for the two distinct scenarios. In both situations, GPT-4 and LLaVA possess the highest BERTScores, demonstrating superior semantic capabilities compared to other models. Additionally, regarding ROUGE, we can see: (1) all models perform better in the home scenario setting; (2) LLaVA displays relatively high ROUGE values and CogVLM excels in a more challenging environment; (3) GPT-4 performs worse than smaller-scaled models. The first observation suggests that the increased visual complexity of the supermarket environment (i.e., a greater variety of objects and more colors in the image) makes the VIALM task more challenging, highlighting the need for models to generalize with in more challenging environments (RQ2). Combining the second and third observations and analyzing the generated guidance, we identify GPT-4’s significant drawback: its output is excessively redundant and lengthy, resulting in low ROUGE scores. In contrast, CogVLM generates the shortest outputs, while LLaVA exhibits the best performance with a moderate guidance length and high ROUGE scores.
@@
Table 5: Human Evaluation Fail Rates. ‘Not Grounded Rate’ is calculated by the sample proportion with Correctness $\leq$ 3 (Score range: 1–5). ‘Not Fine-grained Rate’ is derived from Actionability $\leq$ 3.
4.2 Human Evaluation Results
Figure 4 (b) presents the human evaluation results. To mitigate bias, two independent participants conducted evaluations. In each scenario, 50 random samples were selected. Pearson and Spearman correlation coefficients between the two annotators for each dimension range from 0.7 to 0.9, indicating a strong consensus on the models’ performance. First, it is once again observed that all models provide more accurate answers in home settings than in supermarket settings. Second, as shown in Figure 4 (b), GPT-4 excels in performance across all three dimensions, with CogVLM and LLaVA closely following in second place. Subsequently, we analyze the results for the three specific dimensions:
$\bullet$ In Correctness, GPT-4 performs the best, with CogVLM slightly surpassing LLaVA. It indicates that CogVLM understands the visual environment better than LLaVA, and therefore leads to a superior environment grounding ability.
[…]
@@ 4.2 Human Evaluation Results
This analysis reveals two limitations shared by all models, summarized as the answer to RQ2: ‘not grounded’ (indicated by low correctness) and ‘not fine-grained’ (indicated by low actionability). Table 5 presents the failure rates. It is notable that, despite GPT-4’s superior performance, 25.7% of its responses are not grounded and 32.1% not fine-grained. Additionally, only GPT-4 with a very large scale is capable of providing some general tactile descriptions, while the other five smaller models largely lack this ability.
Case Study.
@@
To better understand the above results, we present an example case in Figure 5 where the question seeks guidance to obtain mustard at a supermarket. The predictions of GPT-4, CogVLM, and LLaVA yield the following observations: (1) GPT-4 generates long output with general guidance but lacks image grounding. (2) CogVLM exhibits better environment grounding, correctly specifying the object’s position, yet it lacks detailed step-by-step guidance. (3) LLaVA’s output emphasizes guiding the user to the specific object but appears to have limited understanding of the environment. (4) GPT-4 generates general tactile guidance, whereas CogVLM and LLaVA entirely ignore tactile descriptions.
@@ 4.3 Discussions
Grounded Fine-grained
Model RG_L BERT_S RG_L BERT_S
Supermarket
GPT-4 0.072 0.245 0.115 0.588
CogVLM 0.144 0.14 0.172 0.495
MiniGPT 0.102 0.202 0.136 0.485
Qwen-VL 0.09 0.169 0.138 0.495
LLaVA 0.097 0.26 0.145 0.554
BLIVA 0.062 0.213 0.107 0.463
@@
GPT-4 0.104 0.405 0.166 0.573
CogVLM 0.158 0.284 0.189 0.443
MiniGPT 0.139 0.415 0.177 0.455
Qwen-VL 0.132 0.311 0.193 0.506
LLaVA 0.123 0.404 0.202 0.544
BLIVA 0.091 0.375 0.159 0.514
@@
[…]
It is observed that CogVLM performs well in environment grounding and LLaVA excels in generating fine-grained outputs. CogVLM’s success in environmental grounding might stem from two factors: (1) the incorporation of the more recent Eva-Clip ViT model, enhancing visual perception, and (2) the inclusion of a Visual Expert module in CogVLM’s LLM architecture, augmenting visual-language alignment. LLAVA’s effectiveness in producing fine-grained guidance might come from its advanced LLM component, LLaMA-2.
Enhancing environment grounding demands advanced visual capabilities, and generating fine-grained guidance necessitates better language abilities. Existing models are limited in synergizing both vision and language skills. It requires future efforts to develop LMs with environment-grounded multimodal abilities for providing fine-grained guidance in VIA.
@@ 5 Conclusion
In this study, we defined the task of VIALM and carried out a comprehensive survey of LMs for VIA, encompassing LLMS, VLMs, and embodied agents. Furthermore, we constructed a benchmark dataset with comprehensive coverage and high difficulty. Following this, we conducted a benchmark evaluation to measure the zero-shot performance of LMs, aiming to examine the potential and limitations of current SOTA models in the VIALM task. The experimental results revealed two primary limitations in the models: their ability to generate grounded and fine-grained guidance.